Bob fenced in a rectangular garden in his yard. The length of the rectangle is 6 feet longer than the width and the perimeter is 104 feet. What is the width of the rectangle?
Answer by guitarguy23(2)

Honestly, im not how to solve it correctly. But heres what i do have.
Go ahead and divide 104 by 4 (sides). You get 26
Since the legnth is 6 feet longer, just add 6 to the legnth and subtract 6 from the width, you end up with the legnth of both sides being 32, and the width of both sides being 20
Check:
Perimiter is adding the 4 sides together.
20+20=40
32+32=64
64+40=104